# Superallowed $0^{+}\rightarrow 0^{+}$ $\beta$ decay studies at GANIL and upcoming opportunities with DESIR and S3-LEB

### v1

@@ 2 Experimental program
The other facility for performing SA measurements at GANIL is DESIR, which will receive beams from both SPIRAL1 and S3-LEB. At DESIR, these beams will be purified first by a High Resolution Separator (HRS) [41] which is designed to have a mass resolution of $m/\Delta m=2\times 10^{4}$. Further improvement in the beam quality will be done by cooling and bunching the beam in a gas filled Paul trap called General Purpose Ion Buncher (GPIB) [42] which can be coupled to a double Penning trap called PIPERADE (PIèges de PEnning pour les RAdionucléides à DEsir) [43]. PIPERADE is currently being commissioned and offline tests have demonstrated a mass resolution of $m/\Delta m=10^{5}$ [43]. Since the beam purification at DESIR does not involve a gas cell or laser ionization, nuclei beyond 54Ni (see Fig. 2) can be studied at DESIR with beams from SPIRAL1 obtained using the TULIP target ion source [25, 26]. The highly purified beam thus obtained will then be sent to different experimental setups placed in the DESIR hall, a layout of which is shown in Fig. 5.
@@
As part of DESIR’s BESTIOL program, a dedicated decay spectroscopy station equipped with a moving tape drive system, charged particle detectors, and gamma detectors will be provided for decay studies. Additionally there will be provisions for users to set up their own detection systems that can be coupled to a mobile tape transport system. Within the DETRAP program, DESIR will host high-resolution mass separation and measurement devices, such as the HRS [41], PIPERADE [43], and, in the near future, a Multi-Reflection Time-of-Flight Mass Spectrometer (MR-TOF-MS) based on the PILGRIM design [35]. These devices will have the capability to perform mass measurements with the precision required for SA $\beta$ decay measurements. In addition, through its LUMIERE program, DESIR will also offer opportunities to perform laser spectroscopy of RIBs or conduct $\beta$-decay studies on laser-polarized RIBs. With multiple setups DESIR will serve as a low-energy, high-precision facility, where all the SA observables can be measured with the same primary radioactive beam.
@@ 3 Concluding remarks
While in the current article we limited our discussions to SA $0^{+}\rightarrow 0^{+}$ Fermi $\beta$ decays, a similar effort is ongoing within the French research community to determine $V_{ud}$ from SA mirror decays [44]. Beyond CKM unitarity tests, the search for BSM physics also includes looking for the existence of scalar or tensor currents, which go beyond the vector-axial vector (V-A) nature of weak interactions. The search for such exotic currents is carried out by precisely measuring the $\beta$ energy spectrum or by performing correlation measurements. At GANIL, within the framework of the b-STILED (b: improved Search for Tensor Interactions in nucLear bEta Decay) project, experiments were conducted at LISE and SPIRAL1 to study the $\beta$ energy spectrum in the decay of 6He [45]. Matter’s Origin from RadioActivity (MORA) is another high-precision experiment that will measure the D correlation in the decay of trapped and laser-polarized 23Mg+ and 39Ca+ ions. The MORA experiment [46] is searching for CP-violating interactions in the V-A framework of the Standard Model and will perform its Phase II measurements at DESIR within the DETRAP program. The ASGARD (Aluminium Superconducting Grid Assembly for Radiation Detection) project [47] will measure the recoil energy of the daughter nucleus to determine the $\beta-\nu$ angular correlation, which is required to determine $V_{ud}$ from SA mirror $\beta$ decays. ASGARD will conduct these experiments at DESIR (and other ISOL facilities) and will also search for scalar and tensor currents by comparing electron capture and $\beta^{+}$ decay rates. WISArD (Weak Interaction Studies with 32Ar Decay) is another high-precision experiment investigating exotic scalar and tensor currents by studying the $\beta-\nu$ angular correlation of 32Ar [48, 49]. The experiment focuses on detecting beta-delayed protons from the daughter nucleus 32Cl. Additionally, they plan to constrain the Fierz interference term by measuring beta-decay spectrum shape. While WISArD is currently setup at ISOLDE, the collaboration plans to replicate the setup to perform similar measurements on 20Mg at DESIR [49].

### v2

@@ 2.2 Upcoming superallowed beta decay studies at SPIRAL1 and LISE at GANIL
The other facility for performing SA measurements at GANIL is DESIR, which will receive beams from both SPIRAL1 and SPIRAL2 via S3-LEB. At DESIR, these beams will be purified first by a High Resolution Separator (HRS) [54] which is designed to have a mass resolution of $m/\Delta m=2\times 10^{4}$. Further improvement in the beam quality will be done by cooling and bunching the beam in a gas filled Paul trap called General Purpose Ion Buncher (GPIB) [55] which can be coupled to a double Penning trap called PIPERADE (PIèges de PEnning pour les RAdionucléides à DEsir) [56]. PIPERADE is currently being commissioned and offline tests have demonstrated a mass resolution of $m/\Delta m=10^{5}$ [56]. The ultimate aim is to attain a mass resolution of $m/\Delta m=10^{7}$. Since the beam purification at DESIR does not involve a gas cell or laser ionization, non-refractory nuclei beyond 54Ni (see Fig. 2) can be studied at DESIR with beams from SPIRAL1 obtained using the TULIP target ion source [38, 39]. The highly purified beam thus obtained will then be sent to different experimental setups placed in the DESIR hall, a layout of which is shown in Fig. 5.
@@
As part of DESIR’s BESTIOL program, a dedicated decay spectroscopy station equipped with a moving tape drive system, charged particle detectors, and gamma detectors will be provided for decay studies. Additionally there will be provisions for users to set up their own detection systems that can be coupled to a mobile tape transport system. The availability of high purity beams, and experimental setups similar to those described in sections 2.1 and 2.2, make DESIR an excellent facility for studying the half-life and branching ratio of superallowed beta decays.
Furthermore, within the DETRAP program, DESIR will host high-resolution mass separation and measurement devices, such as the HRS [54], PIPERADE [56], and, in the near future, a Multi-Reflection Time-of-Flight Mass Spectrometer (MR-TOF-MS) based on the PILGRIM design [48]. These devices will have the capability to perform mass measurements with the precision required for SA $\beta$ decay measurements. In addition, through its LUMIERE program, DESIR will also offer opportunities to perform laser spectroscopy of RIBs or conduct $\beta$-decay studies on laser-polarized RIBs. With multiple setups DESIR will serve as a low-energy, high-precision facility, where all the SA observables can be measured with the same primary radioactive beam.
@@ 3 Concluding remarks
While in the current article we limited our discussions to SA $0^{+}\rightarrow 0^{+}$ Fermi $\beta$ decays, a similar effort is ongoing within the French research community to determine $V_{ud}$ from SA mirror decays [57]. Beyond CKM unitarity tests, the search for BSM physics also includes looking for the existence of scalar or tensor currents, which go beyond the vector-axial vector (V-A) nature of weak interactions. The search for such exotic currents is carried out by precisely measuring the $\beta$ energy spectrum or by performing correlation measurements. At GANIL, within the framework of the b-STILED (b: improved Search for Tensor Interactions in nucLear bEta Decay) project, experiments were conducted at LISE and SPIRAL1 to study the $\beta$ energy spectrum in the decay of 6He [58]. Matter’s Origin from RadioActivity (MORA) is another high-precision experiment that will measure the D correlation in the decay of trapped and laser-polarized 23Mg+ and 39Ca+ ions. The MORA experiment [59, 60] is searching for CP-violating interactions in the V-A framework of the Standard Model and will perform its Phase II measurements at DESIR within the DETRAP program. The ASGARD (Aluminium Superconducting Grid Assembly for Radiation Detection) project [61] will measure the recoil energy of the daughter nucleus to determine the $\beta-\nu$ angular correlation, which is required to determine $V_{ud}$ from SA mirror $\beta$ decays. ASGARD will conduct these experiments at DESIR (and other ISOL facilities) and will also search for scalar and tensor currents by comparing electron capture and $\beta^{+}$ decay rates. WISArD (Weak Interaction Studies with 32Ar Decay) is another high-precision experiment investigating exotic scalar and tensor currents by studying the $\beta-\nu$ angular correlation of 32Ar [62, 63]. The experiment focuses on detecting beta-delayed protons from the daughter nucleus 32Cl. Additionally, they plan to constrain the Fierz interference term by measuring beta-decay spectrum shape. While WISArD is currently setup at ISOLDE, the collaboration plans to replicate the setup to perform similar measurements on 20Mg at DESIR [63].